# Large Scale Multi-Task Bayesian Optimization with Large Language Models

### v1

@@ Abstract
In multi-task Bayesian optimization, the goal is to leverage experience from optimizing existing tasks to improve the efficiency of optimizing new ones. While approaches using multi-task Gaussian processes or deep kernel transfer exist, the performance improvement is marginal when scaling to more than a moderate number of tasks. We introduce a novel approach leveraging large language models (LLMs) to learn from, and improve upon, previous optimization trajectories, scaling to approximately 2000 distinct tasks. Specifically, we propose an iterative framework in which an LLM is fine-tuned using the high quality solutions produced by BayesOpt to generate improved initializations that accelerate convergence for future optimization tasks based on previous search trajectories. We evaluate our method on two distinct domains: database query optimization and antimicrobial peptide design. Results demonstrate that our approach creates a positive feedback loop, where the LLM’s generated initializations gradually improve, leading to better optimization performance. As this feedback loop continues, we find that the LLM is eventually able to generate solutions to new tasks in just a few shots that are better than the solutions produced by “from scratch” by Bayesian optimization while simultaneously requiring significantly fewer oracle calls.
Machine Learning, ICML
@@ 1 Introduction
Multi-task optimization seeks to use related, previously solved tasks to accelerate the optimization of new ones. Multi-task optimization appears naturally in a variety of domains where similar problems are encountered repeatedly, such as hyperparameter optimization, material science, database query optimization, and drug design. Formally, suppose we have tasks $\{1,2,\ldots,T\}$, each associated with its own objective function $f_{t}(\mathbf{x})$. In multi-task Bayesian optimization (BO), we seek to find some $\mathbf{x}^{*}_{t}$ such that
$\mathbf{x}_{t}^{*}=\underset{\mathbf{x}\in\mathcal{X}}{\arg\min}f_{t}(\mathbf{x}),$ (1)
for each task $t$. We focus on the setting where, for each task, we have collected a dataset $D_{t}$ of observations, and we wish to leverage this data when optimizing unseen test tasks.
Multi-task BO methods have predominantly used Gaussian processes (GPs) and/or a variety of shared weight neural network feature extractors to jointly learn correlations across tasks (Swersky et al., 2013; Perrone et al., 2018; Patacchiola et al., 2020; Hakhamaneshi et al., 2022). A standard approach involves placing a multi-output GP over the input-task space, decomposing the kernel as an input kernel $k(\mathbf{x},\mathbf{x}^{\prime})$ and a task kernel $k(t,t^{\prime})$. Despite their effectiveness, many of these methods—with the notable exception of recent work such as Wang et al. (2024c)—tend to saturate in performance after tens of training tasks and do not extract additional optimization performance on new tasks when given hundreds or thousands of related tasks for training.
@@
This approach creates a self-reinforcing feedback loop: BO generates high-quality solutions that we can leverage to fine-tune the LLM; the fine-tuned LLM, in turn, produces better initializations that improve BO performance. Over time, the LLM learns to directly generate solutions that are highly competitive, enabling top-of-$k$-samples from the LLM (requiring just a few oracle calls) to outperform full “from scratch” BO runs (requiring a large number of oracle calls). This iterative improvement enables BOLT to scale to and still extract value from thousands of tasks.
We validate BOLT on two diverse and challenging domains where many related optimization tasks are available.
Application 1: Database query plan optimization. In database query plan optimization, the goal is to optimize execution plans in PostgreSQL to minimize runtime. Efficient query planning is critical for database performance, and traditional optimizers such as PostgreSQL’s built-in query planner rely on hand-crafted heuristics and cost models. We show that an LLM fine-tuned using BOLT can generate superior plans, sometimes outperforming PostgreSQL’s own optimizer in a single-shot setting.
Application 2: Antimicrobial peptide (AMP) design. Antimicrobial peptides are small peptides that kill (pathogenic) bacteria. We use a software package, APEX (Wan et al., 2024), which can predict the minimum inhibitory concentration (MIC) of peptides against various pathogens.
In this setting, each task is a distinct extinct peptide used as a template, and the goal is to enhance its antimicrobial activity through a limited number edits. We demonstrate that our approach continues to improve as more tasks are introduced and, in the few-shot setting, can eventually generate peptides that outperform those found through full “from-scratch” BO runs.
@@
    We validate BOLT through experiments on two challenging domains: database query optimization and antimicrobial peptide design. Results demonstrate that initializations generated by BOLT continue to improve performance as the number of tasks grows.
@@ 2 Background
Bayesian Optimization (BO) (Močkus, 1975; Snoek et al., 2012) is an iterative approach to optimize black-box functions in a sample-efficient manner. On each step of the optimization, a supervised probabilistic surrogate model (usually a Gaussian Process (GP) (Rasmussen, 2003)) is conditioned on all data collected so far. Then, the surrogate model’s predictive posterior distribution $p(y\mid\mathbf{x},D)$ is used to decide what data point(s) should be evaluated next, typically by maximizing some acquisition function, defined with respect to $p(y\mid\mathbf{x},D)$, which guides the exploration-exploitation trade off. Finally, the selected data points are evaluated on the black-box function and added to the dataset. This iterative process continues until the evaluation budget is reached.
@@
Structured optimization via latent space BO. BO has recently been applied to optimizing structured search spaces, such as molecular and amino acid sequences, by leveraging latent space Bayesian optimization. This approach incorporates a variational autoencoder (VAE) to map structured inputs into a continuous latent space, where BO is performed (Eissman et al., 2018; Tripp et al., 2020; Grosnit et al., 2021; Siivola et al., 2021; Stanton et al., 2022; Maus et al., 2022). Structured inputs $\mathbf{x}$ (e.g., amino acid sequences) are mapped to continuous latent representations $\mathbf{z}$ by the VAE encoder $\Phi(\mathbf{x})$. This creates a transformed continuous (latent) representation of the structured search space where BO can be directly applied. The corresponding latent candidate points are then decoded by the VAE decoder, $\Gamma(\mathbf{z})$, to reconstruct structured outputs for evaluation. For large combinatorial structured search spaces, such as the space of organic molecules or the space of all peptide amino acid sequences, the latent space of the VAE is typically high-dimensional (on the order of several hundred dimensions) in order to represent the large structured space effectively.
@@
Optimizing antimicrobial peptides. In antimicrobial peptide design, we seek peptides (sequences of amino acids) that minimize the MIC (minimum inhibitory concentration, measured in $\mu$ mol $\text{L}^{-1}$) for some target bacterial pathogen. MIC is a measure of the concentration of the peptide required to inhibit growth of the target bacterial pathogen (Kowalska-Krochmal & Dudek-Wicher, 2021). A key challenge in antimicrobial peptide design is that many modern bacterial pathogens have developed resistance to modern antibiotics. To solve this challenge, Wan et al. (2024) propose designing new peptides with high sequence similarity to template peptides mined from extinct organisms. The template peptides themselves do not typically achieve sufficiently low MIC for target bacteria pathogens. However, since these template peptides have not been encountered in nature for thousands for years, modern antimicrobial resistant bacteria have not evolved resistance to them. It follows that new peptides are more likely to evade antibiotic resistance if they are designed to be similar to the extinct template sequences. We employ this strategy, optimizing antimicrobial peptides with a minimum threshold sequence similarity to the extinct template peptides from Wan et al. (2024). We also employ latent space BO to optimize over the structured, combinatorial space of amino acid sequences.
@@
Optimizing database query plans. Query optimization in data management systems involves translating a declarative SQL query into an execution plan that efficiently retrieves the correct results (Graefe & McKenna, 1993). This problem has been extensively researched in the field of data management (Selinger et al., 1989), as the difference in execution time between an optimal and a poorly chosen query plan can be several orders of magnitude (Leis et al., 2015). Since individual query plans are composed of discrete characteristics (e.g. join order trees), the search space of possible query plans is structured and combinatorial. We therefore employ latent space BO. We use the string representation for query plans proposed by Tao et al. (2025) to pre-train a VAE model that maps the structured space of query plans to a continuous latent space where BO can be applied.
@@
Database query plan optimization with right-censored observations. In database query optimization, our black-box objective function measures the execution latency of the query plan. The latency difference between a “good” query plan and a “bad” one can be orders of magnitude (Leis et al., 2015). This can lead to the majority of optimization runtime being taken up by evaluating a small number of poorly performing plans. A natural solution to this problem is to timeout objective function evaluations after they have reached some threshold latency $\tau$, resulting in right-censored observations. A right-censored observation is an observation at data point $\mathbf{x}$ where we observe only that $y\geq\tau$ for some chosen timeout threshold $\tau$, rather than observing the typical noisy objective value $y$. Prior work has been done to extend Bayesian optimization methods to the setting of right-censored observations. Hutter et al. (2013); Eggensperger et al. (2020) extended Bayesian optimization methods to the setting of right-censored observations by introducing an EM-like algorithm to impute the values of censored observations. Eggensperger et al. (2020) expanded on this, defining a single surrogate model capable of being conditioned on the combination of censored and uncensored data gathered during optimization.
@@
Tao et al. (2025) extend this to the setting of approximate GP surrogate models. Since we focus on tasks that involve large function evaluation budgets, we employ Tao et al. (2025)’s proposed method of modeling censored data with approximate GPs.
@@ 3 Bayesian Optimization with LLM Transfer (BOLT)
We propose an iterative framework for using large language models (LLMs) to improve Bayesian optimization (BO) performance across a family of related tasks. We are given a set of $T$ training tasks defined by objective functions $f_{1}(\mathbf{x}),...,f_{T}(\mathbf{x})$. We additionally assume that, for each objective function we have a context or task description $C[f_{t}]$ that can be a natural language or other input description that differentiates $f_{t}$ from any other task in the application domain. For example, this might be a SQL query we are trying to optimize.
For each training task, we assume we have optimized the objective with some BO procedure, resulting in the optimization trajectories $\{\mathcal{D}^{\star}_{t}\}_{t=1}^{T}$, with each $\mathcal{D}^{\star}_{t}$ containing the top-$K$ observations from the trajectory for the $t^{\text{th}}$ task.
Our goal is to leverage this training data to learn an LLM-based “initialization policy” $\pi$ that, when presented with new related tasks $\left\{f_{T+1}(\mathbf{x}),C[f_{T+1}(\mathbf{x})]\right\}$, proposes a high-quality set of candidate solutions for BO to further refine.
@@
Because the LLM and BO only interact through generating initialization and generating fine tuning data respectively, our approach here is relatively agnostic to the specific underlying implementation of BO used to optimize for each task. This enables the straightforward use of the full range of recent BO advances on multi-task, constrained, and other optimization settings.
3.1 Initializing BOLT.
At initialization for a workload of tasks, we have only an un-tuned LLM BOLT-00 that is generally useless for the task setting because it is unaware of even the specific format for candidate suggestions. For the first iteration, we solve $T$ optimization tasks with a single-task Bayesian optimization routine where we initialize BO using some standard initialization procedure. We run optimization on each of the $T$ initial tasks, and extract the optimization trajectories $\{\mathcal{D}^{*}_{i}\}_{i=1}^{T}$ from each run.
The number of tasks $T$ that we collect at initialization time–as well as during each round of the BOLT “outer-loop” represents a non-trivial trade-off due to the computational cost of both running BO and the cost of fine-tuning the LLM. Fine-tuning the LLM more frequently results in both additional computational and monetary costs, but allows subsequent BO runs to complete more efficiently.
In this paper, we generally erred on the side of lower monetary cost in exchange for additional compute cost. Specifically, while processing an entire workload of training and test tasks, we only fine-tune the LLM three times.
3.2 LLM Fine-Tuning
The fine-tuning process employs supervised learning using OpenAI’s GPT-4-mini-0718 model through their API. From the optimization trajectories $\{\mathcal{D}^{*}_{i}\}_{i=1}^{T}$, we extract the top-$K$ observations from each of the $T$ runs completed so far. We use these observations along with the task contexts $\left\{C[f_{t}]\right\}_{t=1}^{T}$ to construct a fine-tuning data $\mathcal{D_{\textrm{ft}}}$. Each training instance contains three components:
@@ 3.2 LLM Fine-Tuning
    A user prompt containing the task-specific context $C[f_{t}]$ (e.g., the SQL query requiring optimization).
@@
{minted} [breaklines, fontsize=]latex System: You are a helpful assistant that provides efficient join orderings for given queries. User: SQL query to be optimized Assistant: Optimized query plan
We fine-tune using OpenAI’s standard fine-tuning API (OpenAI et al., 2024). Specifically, we format our data into the required JSONL format (i.e., prompt-solution pairs) and then upload it via the fine-tuning API to initiate training. The model is trained to minimize the negative log-likelihood of the solution tokens x given the task context $C$:
@@
We note that our approach leverages full model fine-tuning rather than extensive prompt engineering (Lester et al., 2021; Li & Liang, 2021). This allows the model to learn the task requirements through the context-solution pairs in $\mathcal{D_{\textrm{ft}}}$, rather than explicit instructions. However, for scenarios requiring few-shot learning on untrained models, more careful prompt engineering may be beneficial.
@@
3.3 Using the LLM for Multi-Task BO
Once we have a fine‐tuned model, BOLT-$T$, we can leverage the fine-tuned LLM’s capabilities to generate higher-quality initialization points for subsequent optimization tasks.
@@ 3.3 Using the LLM for Multi-Task BO
For a set of $n$ new tasks $\{t_{i}\}_{i=T+1}^{T+n}$, we sample from BOLT-$T$ to generate the same number of initialization points used by the baseline “from scratch” approach. The sampling prompt maintains the same structure as the training prompt without the assistant response. The BOLT-$T$ generated solutions are refined with a standard BO routine, and the top‐$K$ performing solutions for each task $t$—along with their contexts $C[f_{t}]$—are incorporated into the training set for the next round of fine‐tuning.
@@
Closed-loop feedback. This iterative process creates a positive feedback loop between BO and the LLM. After completing each iteration of BO, we augment the LLM fine-tuning dataset with the newly collected high-scoring solutions and their corresponding contexts. Re-training the LLM on this expanded dataset allows it to internalize additional insights over the tasks and produce more effective initialization for the next batch of tasks. Through multiple iterations, the LLM learns to propose increasingly high-quality starting points, thereby jump-starting subsequent BO runs and reducing the time needed to find near‐optimal solutions. Ultimately, this positive feedback loop continually refines the model’s understanding of what a good solution should look like under each context, improving its performance across the entire family of tasks.
Algorithm 1 Inner Loop: LLM-Initialized Bayesian Optimization
0:  Task $t$, context $C[f_{t}]$, fine-tuned LLM $\pi_{n}$, evaluation budget $B$, batch size $b$
0:  Optimized solutions $X^{*}_{t}$
1:  $X_{\text{init}}\leftarrow\pi_{n}(C[f_{t}])$LLM proposes initial candidates}
2:  Evaluate $y_{\text{init}}\leftarrow f_{t}(X_{\text{init}})$
3:  Initial dataset $\mathcal{D}\leftarrow\left(X_{\text{init}},y_{\text{init}}\right)$
4:  Initialize GP surrogate model $\mathcal{GP}(X_{\text{init}},y_{\text{init}})$
5:  for step $i=1$ to $\lfloor B/b\rfloor$ do
6:     Select $X_{\text{next}}\leftarrow\operatorname*{arg\,max}_{\alpha}(x;\mathcal{GP})$Acquire new candidates}
7:     Evaluate $y_{\text{next}}\leftarrow f_{t}(X_{\text{next}})$
8:     $\mathcal{D}\leftarrow\left(X\cup X_{next},y\cup y_{next}\right)$Update dataset}
9:     Update $\mathcal{GP}$ with new observations
10:  end for
11:  Return $X^{*}_{t}\leftarrow\text{top-}K(X)$Best solutions for LLM fine-tuning}
Algorithm 2 Outer Loop: LLM Fine-Tuning via BO Trajectories
0:  Initial dataset $\mathcal{D}_{0}=\{(C[f_{t}],x_{i},y_{i})\}$, initial LLM $\pi_{0}$, iteration count $T$
0:  Fine-tuned LLM $\pi_{T}$
1:  Initialize $\mathcal{D}\leftarrow\mathcal{D}_{0}$, $\pi\leftarrow\pi_{0}$
2:  for iteration $k=1$ to $T$ do
3:     for each task $t$ in batch do
4:        $X^{*}_{t_{j}}\leftarrow$ InnerLoop($t$, $\pi_{k}$, $B$, $b$) {BO with LLM initialization}
5:        $\mathcal{D}\leftarrow\mathcal{D}\cup\{(C[f_{t}],x,y)\,|\,x\in X^{*}_{t}\}$Augment with top solutions}
6:     end for
7:     Fine-tune $\pi_{k}$ on $\mathcal{D}$ using instruction prompting
8:  end for
3.4 Self-Augmentation
As the fine-tuned LLM enhances few-shot generation with more optimization data, it is worth exploring whether the costly sequential BO processes can be minimized. Thus, we explore “self-improvement” methods to refine the LLM policy without the expense of additional optimization runs (Algorithm 3).
Specifically, once an LLM has been fine-tuned using some of the tasks we set aside for training, we prompt it to generate additional solutions for all available tasks in that problem setting. We then score these solutions using the problem’s oracle and fine-tune the LLM again directly with this labeled self-generated data, in a manner similar to self-play in reinforcement learning or self-instruction in LLM training (Haluptzok et al., 2022; Shypula et al., 2024). By filtering and fine-tuning on its own best outputs, the LLM can iteratively teach itself how to propose better solutions.

### v2

@@ Abstract
In multi-task Bayesian optimization, the goal is to leverage experience from optimizing existing tasks to improve the efficiency of optimizing new ones. While approaches using multi-task Gaussian processes or deep kernel transfer exist, the performance improvement is marginal when scaling beyond a moderate number of tasks. We introduce a novel approach leveraging large language models (LLMs) to learn from, and improve upon, previous optimization trajectories, scaling to approximately 1500 distinct tasks. Specifically, we propose a feedback loop in which an LLM is fine-tuned on the high quality solutions to specific tasks found by Bayesian optimization (BO). This LLM is then used to generate initialization points for future BO searches for new tasks. The trajectories of these new searches provide additional training data for fine-tuning the LLM, completing the loop. We evaluate our method on two distinct domains: database query optimization and antimicrobial peptide design. Results demonstrate that our approach creates a positive feedback loop, where the LLM’s generated initializations gradually improve, leading to better optimization performance. As this feedback loop continues, we find that the LLM is eventually able to generate solutions to new tasks in just a few shots that are better than the solutions produced by “from scratch” by Bayesian optimization while simultaneously requiring significantly fewer oracle calls.
@@ 1 Introduction
Multi-task optimization seeks to use related, previously solved tasks to accelerate the optimization of new ones. Multi-task optimization appears naturally in a variety of domains where similar problems are encountered repeatedly, such as hyperparameter optimization, material science, database query optimization, and drug design. Formally, suppose we have tasks $\{1,2,\ldots,T\}$, each associated with its own objective function $f_{t}(\mathbf{x})$. For each task $t\in\{1,2,\ldots,T\}$, we seek to find some $\mathbf{x}^{*}_{t}$ such that
$\mathbf{x}_{t}^{*}=\underset{\mathbf{x}\in\mathcal{X}}{\arg\min}f_{t}(\mathbf{x}).$ (1)
We focus on the setting where, for each task, we have collected a dataset $D_{t}$ of observations, and we wish to leverage this data when optimizing unseen test tasks.
Multi-task Bayesian optimization (BO) methods have predominantly used Gaussian processes (GPs) and/or a variety of shared weight neural network feature extractors to jointly learn correlations across tasks [40, 31, 30, 10]. A standard approach involves placing a multi-output GP over the input-task space, decomposing the kernel as an input kernel $k(\mathbf{x},\mathbf{x}^{\prime})$ and a task kernel $k(t,t^{\prime})$. Despite their effectiveness, many of these methods — with the notable exception of recent work such as Wang et al. [47] — tend to saturate in performance after tens of training tasks and do not extract additional performance improvement on new tasks when given hundreds or thousands of related tasks.
@@
This approach creates a self-reinforcing feedback loop: BO generates high-quality solutions that we can leverage to fine-tune the LLM; the fine-tuned LLM, in turn, produces better initializations that improve BO performance. Over time, the LLM learns to directly generate solutions that are highly competitive, enabling top-$k$-samples from the LLM (requiring just a few oracle calls) to outperform full “from scratch” BO runs (requiring a large number of oracle calls). This iterative improvement enables BOLT to scale and still extract value from thousands of tasks. We validate BOLT on two diverse and challenging domains where many related tasks are available.
Application 1: Database query plan optimization. In database query plan optimization, the goal is to find query plans for SQL queries that minimize runtime. Efficient query planning is critical for database performance [15], and traditional optimizers, such as PostgreSQL, rely on hand-crafted heuristics and cost models [35]. We show that an LLM fine-tuned using BOLT can generate superior plans, sometimes outperforming PostgreSQL’s own optimizer in a single-shot setting.
Application 2: Antimicrobial peptide (AMP) design. Antimicrobial peptides are small peptides that kill (pathogenic) bacteria. We use a software package, APEX [44], which can predict the minimum inhibitory concentration (MIC) of peptides against various pathogens. In this setting, each task is a distinct extinct peptide used as a template, and the goal is to enhance its antimicrobial activity through a limited number of edits. We demonstrate that our approach continues to improve as more tasks are introduced and, in the few-shot setting, can eventually generate peptides that outperform those found through full “from-scratch” BO runs.
@@
    We validate BOLT through experiments on two challenging domains: database query optimization and antimicrobial peptide design. Experimental results demonstrate that initializations generated by BOLT continue to improve performance as the number of tasks grows.
@@ 2 Background
Bayesian Optimization (BO) [26, 38] is an iterative approach to optimize black-box functions in a sample-efficient manner. On each step of the optimization, a supervised probabilistic surrogate model (usually a Gaussian Process (GP) [33]) is conditioned on all data collected so far. Then, the surrogate model’s predictive posterior distribution $p(y\mid\mathbf{x},D)$ is used to decide what data point(s) should be evaluated next, typically by maximizing some acquisition function, defined with respect to $p(y\mid\mathbf{x},D)$, which guides the exploration-exploitation trade off. Finally, selected points are evaluated on the black-box function and added to the dataset. This iterative process continues until the evaluation budget is reached.
Structured optimization via latent space BO.
@@
BO has recently been applied to optimizing structured search spaces, such as molecular and amino acid sequences, by leveraging latent space Bayesian optimization. This approach incorporates a variational autoencoder (VAE) to map structured inputs into a continuous latent space, where BO is performed [5, 43, 9, 37, 39, 24]. Structured inputs $\mathbf{x}$ (e.g., amino acid sequences) are mapped to continuous latent representations $\mathbf{z}$ by the VAE encoder $\Phi(\mathbf{x})$. This creates a transformed continuous (latent) representation of the structured search space where BO can be directly applied. The corresponding latent candidate points are then decoded by the VAE decoder, $\Gamma(\mathbf{z})$, to reconstruct structured outputs for evaluation. For large combinatorial structured search spaces, such as the space of organic molecules or the space of all peptide amino acid sequences, the latent space of the VAE is typically high-dimensional (on the order of several hundred dimensions) in order to represent the large structured space effectively.
Optimizing antimicrobial peptides.
@@
In antimicrobial peptide design, we seek peptides (sequences of amino acids) that minimize the MIC (minimum inhibitory concentration, measured in $\mu$ mol $\text{L}^{-1}$) for some target bacterial pathogen. MIC is a measure of the concentration of the peptide required to inhibit growth of the target bacterial pathogen [14]. A key challenge in antimicrobial peptide design is that many modern bacterial pathogens have developed resistance to modern antibiotics. To solve this challenge, Wan et al. [44] propose designing new peptides with high sequence similarity to template peptides mined from extinct organisms. The template peptides themselves do not typically achieve sufficiently low MIC for target bacteria pathogens. However, since these template peptides have not been encountered in nature for thousands of years, modern antimicrobial resistant bacteria have not evolved resistance to them. It follows that new peptides are more likely to evade antibiotic resistance if they are designed to be similar to the extinct template sequences. We employ this strategy, optimizing antimicrobial peptides with a minimum threshold sequence similarity to the extinct template peptides from Wan et al. [44]. We also employ latent space BO to optimize over the structured space of amino acid sequences.
Optimizing database query plans.
@@
Query optimization in data management systems involves translating a declarative SQL query into an execution plan that efficiently retrieves the correct results [8]. This problem has been extensively investigated in the field of data management [16], as the difference in execution time between an optimal and a poorly chosen query plan can be several orders of magnitude [15]. Since individual query plans are composed of discrete characteristics (e.g. join order trees), the search space of possible query plans is structured and combinatorial. We therefore employ latent space BO. We use the string representation for query plans proposed by Tao et al. [41] to pre-train a VAE model that maps the structured space of query plans to a continuous latent space where BO can be applied.
Database query plan optimization with right-censored observations.
@@
In database query optimization, our black-box objective function measures the execution latency of the query plan. “Good” and “bad” query plans can have latencies differing by multiple orders of magnitude [15]. This can lead to the majority of optimization runtime being taken up by evaluating a small number of poorly performing plans. A natural solution to this problem is to time out objective function evaluations after they have reached some threshold latency $\tau$, resulting in right-censored observations. A right-censored observation is an observation at data point $\mathbf{x}$ where we observe only that $y\geq\tau$ for some chosen timeout threshold $\tau$, rather than observing the typical noisy objective value $y$. Prior work has been done to extend Bayesian optimization methods to the setting of right-censored observations. Hutter et al. [12], Eggensperger et al. [4] extended Bayesian optimization methods to the setting of right-censored observations by introducing an EM-like algorithm to impute the values of censored observations. Eggensperger et al. [4] expanded on this, defining a single surrogate model capable of being conditioned on the combination of censored and uncensored data gathered.
@@
Tao et al. [41] extend this to the setting of approximate GP surrogate models. Since we focus on tasks that involve large function evaluation budgets, we employ Tao et al. [41]’s proposed method of modeling censored data with approximate GPs.
@@ 3 Bayesian Optimization with LLM Transfer (BOLT)
We propose Bayesian Optimization with LLM Transfer (BOLT), an iterative framework for using large language models (LLMs) to improve Bayesian optimization (BO) performance across a family of related tasks. We are given a set of $T$ training tasks defined by objective functions $f_{1}(\mathbf{x}),...,f_{T}(\mathbf{x})$. We additionally assume that, for each objective function we have a context or task description $C[f_{t}]$ that can be a natural language or other input description that differentiates $f_{t}$ from any other task in the application domain. For example, this might be the text of a SQL query we are trying to optimize.
For each training task, we assume we have optimized the objective with some BO procedure, resulting in the optimization trajectories $\{\mathcal{D}^{\star}_{t}\}_{t=1}^{T}$, with each $\mathcal{D}^{\star}_{t}$ containing the top-$K$ observations from the trajectory for the $t^{\text{th}}$ task. Our goal is to leverage this training data to learn an LLM-based “initialization policy” $\pi$ that, when presented with new related tasks $\left\{f_{T+1}(\mathbf{x}),C[f_{T+1}(\mathbf{x})]\right\}$, proposes a high-quality set of candidate solutions for BO to further refine.
@@
Because the LLM and BO only interact through generating initialization and generating fine-tuning data respectively, our approach here is relatively agnostic to the specific underlying implementation of BO used to optimize each task. This enables the straightforward use of the full range of recent BO advances on high-dimensional, constrained, and other optimization settings.
Require : Task $t$, context $C[f_{t}]$, LLM $\pi_{n}$, budget $B$, batch $b$
Ensure : Optimized solutions $X^{*}_{t}$
$X_{\text{init}}\leftarrow\pi_{n}(C[f_{t}])$
  // LLM proposes candidates
Evaluate $y_{\text{init}}\leftarrow f_{t}(X_{\text{init}})$
$\mathcal{D}\leftarrow(X_{\text{init}},y_{\text{init}})$
Initialize $\mathcal{GP}(X_{\text{init}},y_{\text{init}})$
for step $i=1$ to $\lfloor B/b\rfloor$ do
       $X_{\text{next}}\leftarrow\operatorname*{arg\,max}_{\alpha}(x;\mathcal{GP})$
        // Acquire
       $y_{\text{next}}\leftarrow f_{t}(X_{\text{next}})$
       $\mathcal{D}\leftarrow(X\cup X_{\text{next}},y\cup y_{\text{next}})$
       Update $\mathcal{GP}$ with new observations
Return $X^{*}_{t}\leftarrow\text{top-}K(X)$
  // Best solutions
Algorithm 1 Inner Loop: LLM-Initialized Bayesian Optimization
Require : Dataset $\mathcal{D}_{0}=\{(C[f_{t}],x_{i},y_{i})\}$, LLM $\pi_{0}$, iterations $T$
Ensure : Fine-tuned LLM $\pi_{T}$
Initialize $\mathcal{D}\leftarrow\mathcal{D}_{0}$, $\pi\leftarrow\pi_{0}$
for iteration $k=1$ to $T$ do
       foreach task $t$ in batch do
             $X^{*}_{t}\leftarrow$ InnerLoop($t$, $\pi_{k}$, $B$, $b$)
              // Run BO
             $\mathcal{D}\leftarrow\mathcal{D}\cup\{(C[f_{t}],x,y)\,|\,x\in X^{*}_{t}\}$
              // Augment with top solutions
       Fine-tune $\pi_{k}$ on augmented dataset $\mathcal{D}$
       Update model parameters via instruction prompting
Return $\pi_{T}$
  // Final fine-tuned LLM
Algorithm 2 Outer Loop: LLM Fine-Tuning via BO Trajectories
Initializing BOLT.
At initialization for a workload of tasks, we have only an un-tuned LLM BOLT-00 that is generally useless for the task setting because it is unaware of even the specific format for candidate suggestions. For the first iteration, we solve $T$ optimization tasks with a single-task BO routine where we initialize BO using some standard initialization procedure. We run optimization on each of the $T$ initial tasks, and extract the optimization trajectories $\{\mathcal{D}^{*}_{i}\}_{i=1}^{T}$ from each run.
LLM fine-tuning.
The LLM fine-tuning process employs supervised learning using OpenAI’s GPT-4-mini-0718 model through their API. From the optimization trajectories $\{\mathcal{D}^{*}_{i}\}_{i=1}^{T}$, we extract the top-$K$ observations from each of the $T$ runs completed so far. We use these observations along with the task contexts $\left\{C[f_{t}]\right\}_{t=1}^{T}$ to construct a fine-tuning dataset $\mathcal{D_{\textrm{ft}}}$. Each training instance contains:
@@
    A user prompt with the task-specific context $C[f_{t}]$ (e.g., the SQL query requiring optimization).
@@
We fine-tune using OpenAI’s standard fine-tuning API [29]. Specifically, we format our data into the required JSONL format (i.e., prompt-solution pairs) and then upload it via the fine-tuning API to initiate training. The model is trained to minimize the negative log-likelihood of the solution tokens x given the task context $C$:
@@
We note that our approach leverages full model fine-tuning rather than extensive and/or manual prompt engineering [17, 18]. This allows the model to learn the task requirements through the context-solution pairs in $\mathcal{D_{\textrm{ft}}}$, rather than explicit instructions. However, for scenarios requiring few-shot learning on untrained models, more careful prompt engineering may be beneficial.
@@
LLM fine-tuning frequency.
The number of tasks $T$ that we collect at initialization time and during each round of the BOLT “outer-loop” represents a non-trivial trade-off due to the computational cost of both running BO and the cost of fine-tuning the LLM. Fine-tuning the LLM more frequently results in both additional computational and monetary costs, but allows subsequent BO runs to complete more efficiently (with fewer black-box function evaluations). In this paper, we erred on the side of lower monetary cost in exchange for additional cost in black-box function evaluations. Specifically, we fine-tuned an LLM 4 times for the query plan optimization task and 7 times for the antimicrobial peptide design task as shown in Figure 2
Using the LLM for multi-task BO.
@@
Once we have a fine-tuned model, BOLT-$T$, we can leverage the fine-tuned LLM’s capabilities to generate higher-quality initialization points for subsequent optimization tasks. For a set of $n$ new tasks $\{t_{i}\}_{i=T+1}^{T+n}$, we sample from BOLT-$T$ to generate the same number of initialization points used by the baseline “from scratch” approach. The sampling prompt maintains the same structure as the training prompt without the assistant response. The BOLT-$T$ generated solutions are refined with a standard BO routine, and the top-$K$ performing solutions for each task $t$—along with their contexts $C[f_{t}]$—are incorporated into the training set for the next round of fine-tuning.
@@
Closed-loop feedback. This iterative process creates a positive feedback loop between BO and the LLM. After completing each iteration of BO, we augment the LLM fine-tuning dataset with the newly collected high-scoring solutions and their corresponding contexts. Re-training the LLM on this expanded dataset allows it to internalize additional insights over the tasks and produce more effective initialization for the next batch of tasks. Through multiple iterations, the LLM learns to propose increasingly high-quality starting points, thereby jump-starting subsequent BO runs and reducing the time needed to find near-optimal solutions. Ultimately, this positive feedback loop continually refines the model’s understanding of what a good solution should look like under each context, improving its performance across the entire family of tasks.
Require : Tasks $\mathcal{T}$, LLM $\pi_{\theta}$, iterations $T$, criteria $\mathcal{C}$
Ensure : Fine-tuned LLM $\pi_{\theta+T}$
Sample $\mathcal{D}\leftarrow\mathcal{D}_{0}$, $\pi\leftarrow\pi_{0}$
for iteration $k=1$ to $T$ do
       foreach task $t\in\mathcal{T}$ do
             $X_{init}\sim\pi_{\theta+k}(t)$
              // Generate samples
             $X^{\star}_{init}\leftarrow SelectBest(X_{init},C)$
              // Select best samples
      $\mathcal{D}\leftarrow\mathcal{D}\cup\{(C[f_{t}],x,y)\,|\,x\in X^{*}_{init}\}$
        // Augment dataset
       Fine-tune $\pi_{\theta+k}$ on $\mathcal{D}$
Return $\pi_{T}$
  // Final fine-tuned LLM